# Mechanical memories in solids, from disorder to design

### v1

@@ 1 INTRODUCTION
In this article we first examine memory in amorphous solids, which will serve as a simple case from which we can generalize to other solids, and which can even form a useful bridge to memories in mechanical metamaterials. While shape memory—most notably the recovery of a shape upon heating [5, 6, 3]—has important connections with the examples here, it operates by a distinct mechanism, and so our brief review omits it. Remarkably, the wealth of memories to come is revealed by deformations alone.
@@ 1.1 Connecting memory and material
How is the memory of direction connected with the physics of a granular packing? Figure 1c shows a layer of discs that accomplish the same feat [7, 2]. The discs are photoelastic so that when they are viewed through cross-polarizing filters, they reveal a network of load-bearing contacts within the material [8]. As the outer boundary is rotated clockwise, the grains form a strongly anisotropic state [9], with chains of strong contacts spiraling between the outer and inner walls. If the rotation is stopped and restarted in the same direction, this anisotropy remains and a steady state is quickly resumed, corresponding to the top curve in Fig. 1a. If the rotation proceeds in the opposite direction, the particles must negotiate a transition to a different anisotropic structure. This process requires a finite rotation angle to complete before settling into a steady state; it is the source of the transient in the bottom curve in Fig. 1a. A qualitatively similar response is also observed in concentrated suspensions of solid spheres [10]. This example sets the pattern for the discussions that follow: our understanding of a material’s memory is tied to our understanding of its physics.
@@ 1.3 Reversible plasticity and return-point memory
Reflecting on the above experiment, one may be shocked that there is any orderly response at all, from a jumble of particles that are endlessly moving around one another—that the paths of particles are anything but a chaotic mess is stunning, let alone that they encode some amount of meaningful information. Recent work has traced this orderly response to localized groups of particles that collectively rearrange under loading. Notably, many of these rearrangements are undone when the loading is reversed, so that each particle returns to its previous position. This process is hysteretic as schematized in Fig. 2e. Under cyclic loading, the plastic deformation of an entire amorphous solid with many such regions can be perfectly reversed by the end of each cycle—a response termed “reversible plasticity”.
These behaviors bring to mind the classic Preisach model [15] for hysteresis in response to an imposed field or other driving parameter—here, the strain $\gamma$. The model considers many elementary units or “hysterons,” each with just two possible states, “$+$” and “--”. The $i$th hysteron switches to each state when $\gamma\geq\gamma_{i}^{+}$ or $\gamma\leq\gamma_{i}^{-}$, respectively (Fig. 2e). While in many applications the hysterons are notional and their thresholds are fitting parameters, observations show that in this case, each rearranging region acts as a hysteron [16, 17, 18, 11]. Indeed, by measuring each rearrangement’s $\gamma^{\pm}$ in a cycle at one strain amplitude, one can construct a Preisach model that predicts the system’s response at smaller amplitudes [12]. Hysterons and their collective behaviors are a starting point for understanding the wealth of behaviors to come in this paper.
@@
One crucial property of the Preisach model is that since the thresholds $\gamma_{i}^{\pm}$ never change, there is a fixed sequence in which transitions happen as $\gamma$ is increased, and another fixed sequence as $\gamma$ is decreased. A consequence of this property is that whenever the driving is bounded between two values $\gamma_{\text{min}}$ and $\gamma_{\text{max}}$, the system’s evolution is likewise bounded [19]—a rule that resembles the “no-passing” property of sliding charge-density waves [20]. Preceding addresses Sri’s (correct) comment that these are not equivalent. For example, if the $i$th hysteron is observed in the “--” state when $\gamma$ is at both of these bounds, then it can never be in the “$+$” state as long as $\gamma$ stays bounded. This property implies that as long as $\gamma_{\text{min}}\leq\gamma\leq\gamma_{\text{max}}$, revisiting either bound will put all hysterons into the same state as when that bound was last visited [19]—a behavior called return-point memory [21, 19, 3, 22]. One can simulate an ensemble of hysterons with random $\gamma^{\pm}$, and drive it with protocols in Fig. 2f that mimic experiments, measuring the fraction of hysterons that differ during readout. The results in Fig. 2g qualitatively match experiments, and they are consistent with return-point memory: the training cycles established bounds at ±$\pm\gamma_{0}$, and even though the initial readout cycles with $\gamma_{\text{read}}<\gamma_{0}$ change the system, the cycle with $\gamma_{\text{read}}=\gamma_{0}$ places the system back onto the same trajectory as before, restoring its state.
@@
Return-point behavior is also recursive: for a portion of the time that strain is bounded between $[\gamma_{\text{min}},\gamma_{\text{max}}]$, it can be further bounded by $[\gamma^{\prime}_{\text{min}},\gamma^{\prime}_{\text{max}}]$ nested within that outer pair, and so on [21]. Thus for cyclic driving, the state of the system at $\gamma=0$ depends on the history of nested turning points. As shown in Fig. 2(f, g), applying the sequence of strain amplitudes $\gamma_{0}=4\%,2\%$ will generally yield a different outcome than $2\%$ or $4\%$ alone—whereas the result of $\gamma_{0}=2\%,4\%$ is indistinguishable from 4%. Crucially, to avoid revisiting a turning point and losing the new memory, both new bounds must be inside the previous interval. For example, omitting $\gamma<0$ in Fig. 2f, so that $\gamma^{\prime}_{\text{min}}=\gamma_{\text{min}}$, will fail to write a second memory.
@@
To look for memories of $2\%$ and $4\%$, one sweeps the strain or strain amplitude from small to large values, exiting the nested intervals from smallest to largest. Leaving an interval means revisiting one of its turning points and erasing the history of smaller deformations since that memory was written. This loss of history changes the system’s differential response to further increases in strain, indicating a memory [21, 3]. Return-point memory is the same principle that lets a combination lock store multiple values from the back-and-forth rotation of a single knob [3].
@@ 1.4 Memory capacity and computation
Now that we have touched on two types of memory that appear to be qualitatively different, we look at another way of seeing that their physics is distinct: namely, their capacity for multiple memories. Memory acts as a filter to preserve only certain features of history: a memory of direction retains only the sign of the most recent deformation, while in return-point memory, multiple amplitudes can be stored only in descending order. We can also estimate the number of bits required to encode the information in readout—i.e. the base-2 logarithm of how many distinct histories readout can represent. Here too, differences are apparent: for memory of a direction this measure is independent of system size $N$, while for return-point memory it scales with $\sqrt{N}$ [23, 12] and for computer memory it scales as $N$. Whether it is probed qualitatively or quantitatively, memory capacity is a tool for describing the non-equilibrium nature of solids in this article, and of countless other systems [3].
@@ 2 MEMORY AND ANNEALING
In our description of reversible plasticity and return-point memory, it is evident that many repeated cycles of a deformation have the same net effect as a single cycle. Yet an amorphous solid that has been freshly formed or deformed catastrophically (strains of order 100%) must be sheared for many cycles before it shows reversible plasticity and return-point memory—here, memory capacity itself depends on history. Each cycle tends to make fewer changes than the previous one, until finally, for $\gamma_{0}$ below some empirical threshold $\gamma_{c}\sim 0.1$, there are no net changes at all, while above $\gamma_{c}$, changes continue indefinitely [24, 25, 26, 27, 28, 13, 29, 18, 30, 31, 32]. The existence of the critical strain $\gamma_{c}$ is reminiscent of reversible-irreversible transitions in other memory-forming systems, and in solids it seems connected with the yielding transition by which the material loses rigidity and begins to flow [32, 29, 25, 24]. The transient “mechanical annealing” below $\gamma_{c}$ resembles its thermal counterpart: a molecular glass held at sufficient temperature will explore different configurations and ultimately reach ones that are more stable and long-lived.
@@
Figure 3: Relationships between annealing and memory. (a) Memory of direction in an un-annealed experimental sample. Pair correlation function $g(r,\theta)$ represents the average positions of each particle’s neighbors. Black dots mark peak at each $\theta$. Comparison with black circle shows elongation along $-45^{\circ}$. Ellipse fit to the peaks has semiaxes $b$, $c$ [12]. (b) Anisotropy $b/c$ in a cycle with amplitude 3.5%, after various annealing protocols. Ring-down fully relaxes anisotropy seen in unannealed samples. Unlabeled curves: annealed with $\gamma_{0}=3.5\%$ (dashed purple), and two trials with $\gamma_{0}=4\%$ (solid and dotted blue curves). (c) Ring-down protocol to relax anisotropy in (b): after reset with amplitude 50%, material is annealed by slowly decreasing amplitude. (d) Excerpt of transition graph by Regev et al. [23] showing states (circles) reachable by forward (black arrows) and reverse (red) shear, starting at un-annealed state $O$. Most arrows are faded for clarity. Like-colored states form strongly-connected components featuring many limit cycles (examples shown with un-faded arrows). (e) Protocols for writing multiple memories. Left: writing both memories after ring-down annealing. Right: writing the smaller memory after constant-amplitude annealing. (f) Experimental multiple-memory readouts. Curves corresponding to (e) show effect of annealing envelope at larger $\gamma_{\text{read}}$. Switching order of memories (“RD 3,4”) erases the smaller one. Ring-down without writing yields “RD” curve. Panels a, b, c, e, f adapted from Ref. 12 (CC BY 4.0). Panel d adapted with permission from Ref. 23: Regev, Attia, Dahmen, Sastry, and Mungan, Phys. Rev. E 103, 062614 (2021). Copyright 2021 by the American Physical Society.
@@ 3 GLASSY MEMORIES AND GLASSY MATTER
Figure 4: Simple behaviors of interacting hysterons. (a) Transition graph with nodes showing the states of two hysterons with a cooperative interaction. Red and black edges correspond to forward and reverse shear. The rightmost edge is an avalanche in which hysteron 1 switching to “--” forces its partner to immediately do the same [18]. (b) Two frustrated hysterons can store an amplitude of asymmetric shear. Thresholds are represented at left; for clarity, only the crucial dependence of $\gamma^{-}_{1}$ on hysteron 2 is shown. At right, a minimal strain protocol shows that the hysterons’ states upon returning to $\gamma=0$ depend on amplitude, violating return-point memory. Adapted from [53]. (c) Multiperiodic orbit. Due to frustration, 4 hysterons return to the same state only after 2 periods of driving. Panel b courtesy of Chloe Lindeman. Panel c adapted from Ref. 54 (CC BY).
@@ 4.1 Connecting amorphous solids to crumpled sheets
Figure 5: Memory from hysterons in crumpled and folded sheets. (a) Force-versus-displacement curves as a pre-crumpled sheet is cyclically compressed from two ends. After tracing two nested sub-cycles, the system can reproduce the outermost hysteresis loop, including much of the detailed structure of the force drops. (b) Height differences in a top-down view of a portion of the sheet, as various elements of the sheet “snap-through” between different stable configurations. (c) Bistability of a circular sheet with a single fold. (d) Origami bellows consisting of four bistable units, tracing out one possible path in its $16$-state transition graph. Panels a and b reproduced with permission from Ref. 63 (CC BY-NC-ND 4.0). Panel c adapted with permission from Ref. 64: F. Lechenault and M. Adda-Bedia, Phys. Rev. Lett. 115, 235501 (2015). Copyright 2015 by the American Physical Society. Panel d adapted with permission from Ref. 65 (CC BY 4.0).
As a first example, consider the mechanical response of a cyclically-crumpled sheet [63]. Figure 5a shows a force versus displacement curve from a thin plastic sheet that has first been prepared by crumpling it and then unfurling it into an approximately flat configuration, and then cyclically compressing it by a time-varying displacement $\Delta$. The curves display nested hysteresis loops, with a series of intermittent force drops throughout. In the protocol shown in the inset, a large driving cycle is applied first, followed by two smaller cycles; when the large driving cycle is applied afterwards, a nearly-identical curve results, even reproducing the dominant force drops. These are the experimental hallmarks of return-point memory. Shohat, Hexner, & Lahini [63] were able to find the hysterons responsible for this macroscopic response, in the form of mesoscopic regions that “snap through” between different stable configurations, pictured in Fig. 5b. Although return-point memory does not require any coupling between the hysterons, Ref. 63 teased apart the coupling between pairs of hysterons in their sheet, showing that their interactions can be cooperative or frustrated.
@@ 4.2 Origami hysterons
One way to elucidate memory and transition pathways in a randomly-crumpled sheet is to study hysterons in sheets with just a few folds. Perhaps the simplest case is shown in Fig. 5c: When a point lying on a crease is pushed normal to the sheet, it can snap through to another stable configuration [66]. Adding folds [64] or changing the shape of the boundary can control the switching thresholds [67] of this kind of hysteron. Intriguingly, bistability survives even when a hole is cut through the sheet along the fold [68], showing that bistability does not require a sharp singularity. Changing the size and shape of the hole offers further control of the switching thresholds.
There is another way to achieve bistability: some origami have folding modes that traverse an energy barrier [69, 70, 71, 72]. By connecting several such units together, one can build up metamaterials with many internal states and a set of allowed transitions between them. The bellows in Fig. 5d shows one such design that possesses memory of its previous deformation [65]. Each of its four modules is based on the same folding pattern, but with slightly different angles so that each module expands and collapses at a different force threshold $F^{\pm}_{i}$. Labeling the deployed state as $1$ and the collapsed state as 00, one may thus list the states of the system with four-bit strings such as $0110$. Any of the $2^{4}$ possible states of this system can be reached through the appropriate series of compressions and expansions, as shown by the transition graph in Fig. 5d; the bellows has return-point memory. A similar mechanical behavior can be realized in a flexible truss structure [73], which resembles the origami bellows with faces removed, and edges replaced by springs.
Figure 6: Coupled hysterons in mechanical metamaterials. (a-d) Corrugated silicone sheet with three ridges that act as hysterons upon vertical compression by a distance $U$. (b) The mechanical response of the system shows sharp drops and recoveries of the force upon loading and unloading to different amplitudes, where each hysteron transitions between its states. (c) Transition graph corresponding to panels (a) and (b). (d) The system experiences a proliferation of new states and transitions if the top plate is tilted by a small amount. (e) A metal sheet with a lattice of bistable dimples can access a multitude of mechanically stable configurations; two of them shown here. Panels a-d reproduced with permission from Ref. 74 (CC BY 4.0). Panel e courtesy of Keith Seffen [75].

### v2

@@ 1 INTRODUCTION
In this article we first examine memory in amorphous solids, which will serve as a foundation from which we can generalize to other solids, and which can even form a useful bridge to memories in mechanical metamaterials. While shape memory—most notably the recovery of a shape upon heating [5, 6, 3]—has important connections with the examples here, it operates by distinct mechanisms, and so our brief review omits it. Remarkably, the wealth of memories to come is revealed by deformations alone.
@@ 1.1 Connecting memory and material
Returning to the binary digit in the bucket of sand: how is the memory of direction connected with the physics of a granular packing? Figure 1c shows a layer of discs that accomplish the same feat [7, 2]. The discs are photoelastic so that when they are viewed through cross-polarizing filters, they reveal a network of load-bearing contacts within the material [8]. As the outer boundary is rotated clockwise, the grains form an anisotropic state [9] with chains of strong contacts spiraling between the outer and inner walls. If the rotation is stopped and restarted in the same direction, this anisotropy remains and a steady state is quickly resumed, corresponding to the top curve in Fig. 1a. If the rotation proceeds in the opposite direction, the particles must negotiate a transition to a different anisotropic structure. This process requires a finite rotation angle to complete before settling into a steady state; it is the source of the transient in the bottom curve in Fig. 1a. A qualitatively similar response is also observed in concentrated suspensions of solid spheres [10]. This example sets the pattern for the discussions that follow: our understanding of a material’s memory is tied to our understanding of its physics.
@@ 1.3 Reversible plasticity and return-point memory
Reflecting on the above experiment, one may be shocked that there is any orderly response at all, from a jumble of particles that are endlessly moving around one another. That the paths of particles are anything but a chaotic mess is stunning, let alone that they encode some amount of meaningful information. Recent work has traced this orderly response to localized groups of particles that collectively rearrange under loading (Fig. 2d). Notably, many of these rearrangements are undone when the loading is reversed, so that each particle returns to its previous position. This process is hysteretic as schematized in Fig. 2e. Under cyclic loading, the plastic deformation of an entire amorphous solid with many such regions can be perfectly reversed by the end of each cycle—a response termed “reversible plasticity”.
These behaviors bring to mind the classic Preisach model [15, 16] for hysteresis in response to an imposed field or other driving parameter—here, the strain $\gamma$. The model considers many elementary units or “hysterons,” each with just two possible states, “$+$” and “--”. The $i$th hysteron switches to each state when $\gamma\geq\gamma_{i}^{+}$ or $\gamma\leq\gamma_{i}^{-}$, respectively (Fig. 2e). While in many applications the hysterons are notional and their thresholds are fitting parameters [17, 18, 19], observations show that in this case, each rearranging region acts as a hysteron [20, 21, 22, 11]. Indeed, by measuring each rearrangement’s $\gamma^{\pm}$ in a cycle at one strain amplitude, one can construct a Preisach model that predicts the system’s response at smaller amplitudes [12]. Hysterons and their collective behaviors are a starting point for understanding the wealth of behaviors to come in this review.
@@
One crucial property of the Preisach model is that since the thresholds $\gamma_{i}^{\pm}$ never change, there is a fixed sequence in which transitions happen as $\gamma$ is increased, and another fixed sequence as $\gamma$ is decreased. A consequence of this property is that whenever the driving is bounded between two values $\gamma_{\text{min}}$ and $\gamma_{\text{max}}$, the system’s evolution is likewise bounded [23]—a rule that resembles the “no-passing” property of sliding charge-density waves [24]. For example, if the $i$th hysteron is observed in the “--” state when $\gamma$ is at both of these bounds, then it can never be in the “$+$” state as long as $\gamma$ stays bounded. This property implies that as long as $\gamma_{\text{min}}\leq\gamma\leq\gamma_{\text{max}}$, revisiting either bound will put all hysterons into the same state as when that bound was last visited [23]—a behavior called return-point memory [25, 23, 3, 26]. One can simulate an ensemble of hysterons with random $\gamma^{\pm}$, and drive it with protocols in Fig. 2f that mimic experiments, measuring the fraction of hysterons that differ during readout. The results in Fig. 2g qualitatively match experiments, and they are consistent with return-point memory: the training cycles established bounds at ±$\pm\gamma_{0}$, and even though the initial readout cycles with $\gamma_{\text{read}}<\gamma_{0}$ change the system, the cycle with $\gamma_{\text{read}}=\gamma_{0}$ places the system back onto the same trajectory as before, restoring its state.
@@
Return-point behavior is also recursive: for a portion of the time that strain is bounded between $[\gamma_{\text{min}},\gamma_{\text{max}}]$, it can be further bounded by $[\gamma^{\prime}_{\text{min}},\gamma^{\prime}_{\text{max}}]$ nested within that outer pair, and so on [25]. Thus for cyclic driving, the state of the system at $\gamma=0$ depends on the history of nested turning points. As shown in Fig. 2(f, g), applying the sequence of two strain amplitudes $\gamma_{0}=4\%,2\%$ will generally yield a different outcome than $2\%$ or $4\%$ alone—whereas the result of $\gamma_{0}=2\%,4\%$ is indistinguishable from 4%. Crucially, to avoid revisiting a turning point and losing the new memory, both new bounds must be inside the previous interval. For example, omitting all $\gamma<0$ in Fig. 2f, so that $\gamma^{\prime}_{\text{min}}=\gamma_{\text{min}}$, will fail to write a second memory.
@@
To look for memories of $2\%$ and $4\%$, one sweeps the strain amplitude (or, in some cases, strain) from small to large values, exiting the nested intervals from smallest to largest. Leaving an interval means revisiting one of its turning points and erasing the history of smaller deformations since that memory was written. This loss of history changes the system’s differential response to further increases in strain, indicating a memory [25, 18, 19, 3]. Return-point memory is the same principle that lets a combination lock store multiple values from the back-and-forth rotation of a single knob [3].
@@ 1.4 Memory capacity and computation
Now that we have touched on two types of memory that appear to be qualitatively different, we look at another way of seeing that their physics is distinct: namely, their capacity for multiple memories. Memory acts as a filter to preserve only certain features of history: a memory of direction retains only the sign of the most recent deformation, while in return-point memory, multiple amplitudes can be stored only in descending order. We can also estimate the number of bits required to encode the information in readout—i.e. the base-2 logarithm of how many distinct histories readout can represent. Here too, differences are apparent: for memory of a direction this measure is independent of system size $N$, while for return-point memory it scales with $\sqrt{N}$ [27, 12] and for computer memory it scales as $N$. Whether it is probed qualitatively or quantitatively, memory capacity is a tool for describing the non-equilibrium nature of the solids in this review, and of countless other systems [3].
@@ 2 MEMORY AND ANNEALING
In our description of reversible plasticity and return-point memory, it is evident that many repeated cycles of a deformation have the same net effect as a single cycle. Yet an amorphous solid that has been freshly formed or deformed catastrophically (strains of order 100%) must be sheared for many cycles before it shows reversible plasticity and return-point memory—here, memory capacity itself depends on history. Each cycle tends to make fewer changes than the previous one, until finally, for $\gamma_{0}$ below some empirical threshold $\gamma_{c}\sim 10\%$, there are no net changes at all. In contrast, for $\gamma_{0}>\gamma_{c}$, changes continue indefinitely [28, 29, 30, 31, 32, 13, 33, 22, 34, 35, 36]. The existence of the critical strain $\gamma_{c}$ is reminiscent of reversible-irreversible transitions in other memory-forming systems, and in solids it seems connected with the yielding transition by which the material loses rigidity and begins to flow [36, 33, 29, 28]. The transient “mechanical annealing” below $\gamma_{c}$ resembles its thermal counterpart: a molecular glass held at sufficient temperature will explore different configurations and ultimately reach ones that are more stable and long-lived.
@@
Figure 3: Relationships between annealing and memory. (a) Memory of direction in an un-annealed experimental sample. Pair correlation function $g(r,\theta)$ represents the average positions of each particle’s neighbors. Black dots mark peak at each $\theta$. Comparison with black circle shows elongation along $-45^{\circ}$. Ellipse fit to the peaks has semiaxes $b$, $c$ [12]. (b) Anisotropy $b/c$ in a cycle with amplitude 3.5%, after various annealing protocols. Ring-down fully relaxes anisotropy seen in unannealed samples. Unlabeled curves: annealed with $\gamma_{0}=3.5\%$ (dashed purple), and two trials with $\gamma_{0}=4\%$ (solid and dotted blue curves). (c) Ring-down protocol to relax anisotropy in (b): after reset with amplitude 50%, material is annealed by slowly decreasing amplitude. (d) Excerpt of transition graph by Regev et al. [27] showing states (circles) reachable by forward (black arrows) and reverse (red) shear, starting at un-annealed state $O$. Most arrows are faded for clarity. Like-colored states form strongly-connected components featuring many limit cycles (examples shown with un-faded arrows). (e) Protocols for writing and reading multiple memories. Left: writing both memories after ring-down annealing. Right: writing the smaller memory after constant-amplitude annealing. (f) Experimental multiple-memory readouts. Curves corresponding to (e) show effect of annealing envelope at larger $\gamma_{\text{read}}$. Switching order of memories (“RD 3,4”) erases the smaller one. Ring-down without writing yields “RD” curve. Panels a, b, c, e, f adapted from Ref. 12 (CC BY 4.0). Panel d adapted with permission from Ref. 27: Regev, Attia, Dahmen, Sastry, and Mungan, Phys. Rev. E 103, 062614 (2021). Copyright 2021 by the American Physical Society.
@@ 3 GLASSY MEMORIES AND GLASSY MATTER
Figure 4: Simple behaviors of interacting hysterons. (a) Transition graph with nodes showing the states of two hysterons with a cooperative interaction. Red and black edges correspond to forward and reverse shear. The rightmost edge is an avalanche in which hysteron 1 switching to “--” forces its partner to immediately do the same [22]. (b) Two frustrated hysterons can store an amplitude of asymmetric shear [57]. Thresholds are represented at left; for clarity, only the crucial dependence of $\gamma^{-}_{1}$ on hysteron 2 is shown. At right, a minimal strain protocol shows that the hysterons’ states upon returning to $\gamma=0$ depend on amplitude, violating return-point memory. (c) Multiperiodic orbit. Due to frustration, 4 hysterons return to the same state only after 2 periods of driving. Panel b courtesy of Chloe Lindeman. Panel c adapted from Ref. 58 (CC BY).
@@ 4.1 Connecting amorphous solids to crumpled sheets
Figure 5: Memory from hysterons in crumpled and folded sheets. (a) Force-versus-displacement curves as a pre-crumpled sheet is cyclically compressed from two ends. After tracing two nested sub-cycles, the system can reproduce the outermost hysteresis loop, including much of the detailed structure of the force drops. (b) Height differences in a top-down view of a portion of the sheet, as various elements of the sheet “snap-through” between different stable configurations. (c) Bistability of a circular sheet with a single fold. (d) Origami bellows consisting of four bistable units, tracing out one possible path in its $16$-state transition graph. Panels a and b reproduced with permission from Ref. 67 (CC BY-NC-ND 4.0). Panel c adapted with permission from Ref. 68: F. Lechenault and M. Adda-Bedia, Phys. Rev. Lett. 115, 235501 (2015). Copyright 2015 by the American Physical Society. Panel d adapted with permission from Ref. 69 (CC BY 4.0).
As a first example, consider the mechanical response of a cyclically-crumpled sheet [70, 71, 72, 73, 74, 75, 76, 77, 78]. Figure 5a shows a force versus displacement curve from a thin plastic sheet that has first been prepared by crumpling it and then unfurling it into an approximately flat configuration, and then cyclically compressing it [67]. The curves display nested hysteresis loops, with a series of intermittent force drops throughout. The driving protocol is shown in the inset: A large driving cycle is applied first, followed by two smaller cycles and a final large cycle. Remarkably, the final cycle produces a nearly-identical hysteresis loop as the first, even reproducing the dominant force drops. These are hallmarks of return-point memory. Shohat, Hexner, & Lahini [67] were able to find the hysterons responsible for this macroscopic response: the sheet mesoscopic regions that “snap through” between different stable configurations, pictured in Fig. 5b. Although return-point memory does not require any coupling between the hysterons, Ref. 67 teased apart the coupling between pairs of hysterons in their sheet, showing that their interactions can be cooperative or frustrated. In a subsequent study, Shohat & Lahini [79] identified excess energy dissipation as a harbinger of memory formation in crumpled sheets, forming another parallel between memories in crumpled sheets and in disordered particulate matter.
@@ 4.2 Origami hysterons
One way to elucidate the memory in a randomly-crumpled sheet is to study hysterons in sheets with just a few folds. Perhaps the simplest case is shown in Fig. 5c: when a point lying on a crease is pushed normal to the sheet, it can snap through to another stable configuration [80]. Adding folds [68] or changing the shape of the boundary can control the switching thresholds [81] of this kind of hysteron. Intriguingly, bistability survives even when a hole is cut through the sheet along the fold [82], removing the point-like elastic singularity. Changing the size and shape of the hole offers further control of the switching thresholds.
There is another way to achieve bistability: some origami have folding modes that traverse an energy barrier [83, 84, 85, 86]. By connecting several such units together, one can build up metamaterials with many internal states and a set of allowed transitions between them. The bellows in Fig. 5d shows a design that possesses memory of its previous deformation [69]. Each of its four modules is based on the same folding pattern, but with slightly different angles so that each module expands and collapses at a different force $F^{\pm}_{i}$. Labeling the deployed state as $1$ and the collapsed state as 00, one may thus list the states of the system with four-bit strings such as $0110$. Any of the $2^{4}$ possible states of this system can be reached through the appropriate series of compressions and expansions, as shown by the transition graph in Fig. 5d; the bellows has return-point memory. A similar mechanical behavior can be realized in a flexible truss structure [87], which resembles the origami bellows with faces removed, and edges replaced by springs.
Figure 6: Coupled hysterons in mechanical metamaterials. (a-d) Corrugated silicone sheet with three ridges that act as hysterons upon vertical compression by a distance $U$. (b) The mechanical response of the system shows sharp drops and recoveries of the force upon loading and unloading to different amplitudes, where each hysteron transitions between its states. (c) Transition graph corresponding to panels (a) and (b). (d) The system experiences a proliferation of new states and transitions if the top plate is tilted by a small amount. (e) A metal sheet with a lattice of bistable dimples can access a multitude of mechanically stable configurations; two of them shown here. Panels a-d reproduced with permission from Ref. 88 (CC BY 4.0). Panel e reproduced with permission from Ref. 89.